# Unitary normalizers in finite-dimensional inclusions

### v1

@@ 3.3 Regularity of Inclusions in the General Case
Proof: Fix $k\in\{0,1,\cdots,(s-1)\}$ and let the support of $k^{th}$ row be $Y_{k}=\{x_{1},x_{2},\cdots,x_{r_{k}}\}$. By equation 2.4 and Lemma 3.8, we have
@@
As a consequence, any regular inclusion is isomorphic to tensor product and direct sum of inclusions $M_{n}(\mathbb{C})\subset M_{n}(\mathbb{C}),$ $\mathbb{C}\subset\oplus_{i=0}^{s-1}M_{n}(\mathbb{C})$ and $\mathbb{C}^{n}\subset M_{n}(\mathbb{C})$.
@@
Let $P=P_{i}+P_{k}$ be the corresponding central projection in $\mathcal{A}$. By Proposition 3.8, the induced inclusion
$P\mathcal{B}\subset P\mathcal{A}$
is regular, and moreover
$P\mathcal{N}_{\mathcal{A}}(\mathcal{B})\subset\mathcal{N}_{P\mathcal{A}}(P\mathcal{B})).$
@@
Since the $Q_{j}$ are minimal central projections of $\mathcal{B}$, there exists another minimal projection $Q_{x}$ such that
$UQ_{j}U^{*}=Q_{x}.$
Hence,
@@
Assume $\mathcal{B}$ is commutative. Then the regularity of the inclusion $\mathcal{B}\subset\mathcal{A}$ is equivalent to the inclusion matrix being a normalizer matrix. Indeed, in this case each matrix summand of $\mathcal{B}$ has one-dimensional, that is, $m_{j}=1$ for all $j\in\{0,1,\cdots,r-1\}$. Moreover, the converse of Proposition 3.14 also holds. More precisely, if $\mathcal{B}$ is homogeneous and all entries of the inclusion are equal to $t$, then $\mathcal{B}\subset M_{n}(\mathbb{C})$ is isomorphic to tensor product of regular inclusions
@@ 4.1 Unitary Orthonormal basis
We now turn to the problem of the existence of unitary orthonormal bases for finite-dimensional inclusions, with particular emphasis on bases contained in the normalizer. Building on the combinatorial description of regularity developed in the previous section, we show that the existence of such bases is completely determined by the inclusion matrix, associated dimension data of simple summands and an additional Spectral condition. This characterization reduces the problem to explicit combinatorial and spectral criteria and yields concrete construcions of unitary orthonormal bases in the normalizer. As an application, we obtain an affirmative solution to a finite-dimensional analogue of the Bakshi-Gupta conjecture [3].
@@
$\varphi(\oplus X_{i})=\frac{1}{\sum_{i=0}^{s-1}n_{i}^{2}}\sum_{i=0}^{s-1}n_{i}~\mbox{trace}~(X_{i}),~~X_{i}\in M_{n_{i}},1\leq i\leq(s-1).$
@@
with inclusion matrix
$A^{\prime}=\text{diag}(A_{11},A_{22},\cdots,A_{pp})$
and whose dimension vector of $\tilde{\mathcal{B}}$ is
@@
By dimension counting, we have $n_{k_{i}}=m_{k}r_{k}a_{k_{i}k^{1}}\qquad 1\leq i\leq s_{k}.$ so the above spectral condition becomes
@@
for $1\leq k\leq p$ and satisfies the spectral condition (eq. 4.8).(Note that $\tilde{\varphi}=\sum_{i=1}^{k}\tilde{\varphi}_{i}$)
@@
The following corollary is an analogue of Bakshi-Gupta conjecture [3] for finite dimensional $C^{*}$- algebras.
@@
Proof: If the inclusion admits a unitary orthonormal basis in the normalizer, then it by the definition of basis, the inclusion is regular.
@@
Then the inclusion is $M_{m}\subset\oplus_{i=0}^{s-1}M_{n_{i}}$ whose inclusion matrix is
$A=\begin{bmatrix}a_{0}&a_{1}&\ldots&a_{s-1}\end{bmatrix}^{t}$ (4.9)
Clearly, the above inclusion matrix satisfies the spectral condition and it is a normalizer matrix. So, by the above theorem, it admits a unitary orthonormal basis in the normalizer.
@@
Proof: Let $A=\begin{bmatrix}a_{0}&a_{1}&\cdots&a_{r-1}\end{bmatrix}$ denote the inclusion matrix. Suppose first that $A$ satisfies the spectral condition, then $An^{\prime}=dm^{\prime}$, where $n^{\prime}=\begin{bmatrix}n\end{bmatrix}$ and $m^{\prime}=\begin{bmatrix}1&1&\cdots&1\end{bmatrix}^{t}$. This implies $a_{j}n=d\quad\forall j$. Consequently, all entries of $A$ are equal. Hence $A$ is a normalizer matrix and by Theorem 3.17, the inclusion is regular. Conversely, if the inclusion is regular, then $a_{j}=t\quad\forall j$, for some $t\in\mathbb{N}$. Therefore, $A$ satisfies the spectral condition. This completes the proof. $\Box$
@@
Combining the above two corollaries, suppose that the smaller algebra $\mathcal{B}$ is abelian and that either $\mathcal{A}$ or $\mathcal{B}$ is a matrix algebra. Then the following conditions are equivalent:
@@ 4.2 Depth of an inclusion
where the power of $s\times r$ matrix $A$ is understood us $A^{2}=AA^{t},A^{3}=AA^{t}A$, and so forth. The definition depends only on the equivalence class of $A$ up to permutation.
In parts of the literature, the inclusion matrix is defined as the transpose of the inclusion matrix used in this article. In general, the depths associated with a matrix $A$ and its transpose $A^{t}$ need not coincide. However, when the depth is even, $A$ and $A^{t}$ have the same depth, [5] Theorem 3.16.
@@
If the inclusion $\mathcal{B}\subset\mathcal{A}$ is regular, then the corresponding inclusion matrix $A$ is of depth 2 and the constant $q$ appearing in the definition is bounded from below by square of norm of the inclusion matrix.
@@
$q=\text{max}\{r_{k}\sum_{i=1}^{s_{k}}a_{k_{i}k^{1}}^{2}:\;1\leq k\leq p\}=\lVert A\rVert^{2}.$

### v2

@@ 3.3 Regularity of Inclusions in the General Case
Proof: Fix $k\in\{0,1,\cdots,s-1\}$ and let the support of $k^{th}$ row be $Y_{k}=\{x_{1},x_{2},\cdots,x_{r_{k}}\}$. By equation 2.4 and Lemma 3.8, we have
@@
As a consequence, any regular inclusion is isomorphic to tensor product and direct sum of inclusions $M_{k}(\mathbb{C})\subset M_{k}(\mathbb{C}),$ $\mathbb{C}\subset\oplus_{i=0}^{v-1}M_{l_{i}}(\mathbb{C})$ and $\mathbb{C}^{t}\subset M_{t}(\mathbb{C})$.
@@
Let $P=P_{i}+P_{k}$ be the corresponding central projection in $\mathcal{A}$. By Proposition 3.8, the induced inclusion $P\mathcal{B}\subset P\mathcal{A}$ is regular, and moreover $P\mathcal{N}_{\mathcal{A}}(\mathcal{B})\subset\mathcal{N}_{P\mathcal{A}}(P\mathcal{B})).$
@@
Since the $Q_{j}$ are minimal central projections of $\mathcal{B}$, there exists another minimal projection $Q_{x}$ such that $UQ_{j}U^{*}=Q_{x}.$ Hence,
@@
Assume $\mathcal{B}$ is commutative. Then the regularity of the inclusion $\mathcal{B}\subset\mathcal{A}$ is equivalent to the inclusion matrix being a normalizer matrix. Indeed, in this case each matrix summand of $\mathcal{B}$ has one-dimensional, that is, $m_{j}=1$ for all $j\in\{0,1,\cdots,r-1\}$. Moreover, the converse of Proposition 3.14 also holds. More precisely, if $\mathcal{B}$ is homogeneous and all entries of the inclusion matrix are equal to $t$, then $\mathcal{B}\subset M_{n}(\mathbb{C})$ is isomorphic to tensor product of regular inclusions
@@ 4.1 Unitary Orthonormal basis
We now turn to the problem of the existence of unitary orthonormal bases for finite-dimensional inclusions, with particular emphasis on bases contained in the normalizer. Building on the combinatorial description of regularity developed in the previous section, we show that the existence of such bases is completely determined by the inclusion matrix, associated dimension data of simple summands and an additional spectral condition [1]. While the construction in [1] produces unitary orthonormal bases for several inclusions, these bases need not, in general, be contained in the normalizer. In contrast, our characterization reduces the problem to explicit combinatorial and spectral criteria and yields concrete construcions of unitary orthonormal bases in the normalizer. As an application, we obtain an affirmative solution to a finite-dimensional analogue of the Bakshi-Gupta conjecture [3].
@@
$\varphi(\oplus X_{i})=\frac{1}{\sum_{i=0}^{s-1}n_{i}^{2}}\sum_{i=0}^{s-1}n_{i}~\mbox{trace}~(X_{i}),~~X_{i}\in M_{n_{i}},0\leq i\leq(s-1).$ (4.8)
@@
with inclusion matrix $A^{\prime}=\text{diag}(A_{11},A_{22},\cdots,A_{pp})$ and the dimension vector of $\tilde{\mathcal{B}}$ is
@@
By dimension counting, we have $n_{k_{i}}=m_{k}r_{k}a_{k_{i}k^{1}}\qquad 1\leq i\leq s_{k}$, so the above spectral condition becomes
@@
for $1\leq k\leq p$ and satisfies the spectral condition (eq. 4.9).(Note that $\tilde{\varphi}=\sum_{k=1}^{p}\tilde{\varphi}_{i}$)
@@
The following corollary is an analogue of Bakshi-Gupta conjecture [3] for finite dimensional von Neumann algebras.
@@
Proof: If the inclusion admits a unitary orthonormal basis in the normalizer, then by the definition of basis, the inclusion is regular.
@@
Then the inclusion is $M_{m}\subset\oplus_{i=0}^{s-1}M_{n_{i}}$ whose inclusion matrix is $A=\begin{bmatrix}a_{0}&a_{1}&\ldots&a_{s-1}\end{bmatrix}^{t}$ Clearly, the above inclusion matrix satisfies the spectral condition and it is a normalizer matrix. So, by the above theorem, it admits a unitary orthonormal basis in the normalizer.
@@
Proof: Let $A=\begin{bmatrix}a_{0}&a_{1}&\cdots&a_{r-1}\end{bmatrix}$ denote the inclusion matrix. Suppose first that $A$ satisfies the spectral condition, then $A^{t}n^{\prime}=dm^{\prime}$, where $n^{\prime}=\begin{bmatrix}n\end{bmatrix}$ and $m^{\prime}=\begin{bmatrix}1&1&\cdots&1\end{bmatrix}^{t}$. This implies $a_{j}n=d\quad\forall j$. Consequently, all entries of $A$ are equal. Hence $A$ is a normalizer matrix and by Theorem 3.17, the inclusion is regular. Conversely, if the inclusion is regular, then $a_{j}=t\quad\forall j$, for some $t\in\mathbb{N}$. Therefore, $A$ satisfies the spectral condition. This completes the proof. $\Box$
@@
Combining the above two corollaries, suppose that the smaller algebra $\mathcal{B}$ is commutative and that either $\mathcal{A}$ or $\mathcal{B}$ is a matrix algebra. Then the following conditions are equivalent:
@@ 4.2 Depth of an inclusion
where the power of $s\times r$ matrix $A$ is understood as $A^{2}=AA^{t},A^{3}=AA^{t}A$, and so forth. The definition depends only on the equivalence class of $A$ up to permutation.
In parts of the literature, the inclusion matrix is defined as the transpose of the inclusion matrix used in this article. In general, the depths associated with a matrix $A$ and its transpose $A^{t}$ need not coincide. However, when the depth is even, $A$ and $A^{t}$ have the same depth, see [5] Theorem 3.16.
@@
If the inclusion $\mathcal{B}\subset\mathcal{A}$ is regular, then the corresponding inclusion matrix $A$ is of depth 2 and the constant $q$ appearing in the definition is bounded below by square of norm of the inclusion matrix.
@@
$q=\text{max}\{r_{k}\sum_{i=1}^{s_{k}}a_{k_{i}k^{1}}^{2}:\;1\leq k\leq p\}=\lVert A\rVert^{2},$